OpenStudy (anonymous):

i need help!!here is the question:give the rule for the number sequence.write the net three numbers.1,3,8,19,42,89_,_,_

OpenStudy (anonymous):

1*2+1=3 3*2+2=8 8*2+3=19 19*2+4=42

OpenStudy (anonymous):

Mathe is right

OpenStudy (anonymous):

1x2+1 = 3 3x2+2 = 8 8x2+3 = 19 next 3 numbers = 184, 375, 758